OpenStudy (anonymous):

What are the exact solutions of x2 + 5x + 2 = 0?

OpenStudy (anonymous):

\[x=\frac{-5\pm\sqrt{17}}{2}\]

OpenStudy (anonymous):

\[x = 1/2 (-5-\sqrt(17))\]

OpenStudy (anonymous):

via \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]with \[a=1,b=5,c=2\]

myininaya (myininaya):

\[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

\[x = 1/2 (\sqrt(17)-5)\]

OpenStudy (anonymous):

@myininaya: That's not a solution, that's the quadratic equation!

myininaya (myininaya):

no its the quadratic formula im just showing what this person can use to find the answers

OpenStudy (anonymous):

Quadratic formula and quadratic equation is the same thing.